# Sequence Reconstruction for Single-Deletion Single-Substitution Channel

### v1

@@ II-C The notation $\phi$
Suppose $\bm{x},\bm{x}^{\prime}\in\Sigma_{2}^{8}$ such that
$\bm{x}$ $=01010111$
$~{}\bm{x}^{\prime}$ $=01101011.$
Then we can find
$x_{[n]\backslash 4}$ $=0100111$
$~{}\bm{x}^{\prime}_{[n]\backslash 7}$ $=0110101$
and
$B^{\text{S}}_{1}(x_{[n]\backslash 4},x^{\prime}_{[n]\backslash 7})=\{0100101,0110111\}.$
In fact, we can easily check that $\phi^{4}_{7;x^{\prime}_{6}}(\bm{x})=0100101=\phi^{7}_{3;x_{3}}(\bm{x}^{\prime})$ and $\phi^{4}_{3;x^{\prime}_{3}}(\bm{x})=0110111=\phi^{7}_{6;x_{7}}(\bm{x}^{\prime})$.
@@
For any $j,j^{\prime}\in[n]$ such that $j\leq j^{\prime}$, if $d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=2$, then by Observation 2, we can let $\big{(}S\cap[1,j-1]\big{)}\cup\big{(}T^{L}\cap[j+1,j^{\prime}]\big{)}\cup\big{(}S\cap[j^{\prime}+1,n]\big{)}=\{j_{1},j_{2}\}$. For each $\ell\in\{1,2\}$: if $j_{\ell}\in[1,j-1]\cup[j^{\prime}+1,n]$, let $\bm{z}_{\ell}=\phi^{j}_{j_{\ell};x^{\prime}_{j_{\ell}}}(\bm{x})$ and $\bm{w}_{\ell}=\phi^{j^{\prime}}_{j_{\ell};x_{j_{\ell}}}(\bm{x}^{\prime})$; if $j_{\ell}\in[j+1,j^{\prime}]$, let $\bm{z}_{\ell}=\phi^{j}_{j_{\ell};x^{\prime}_{j_{\ell}-1}}(\bm{x})$ and $\bm{w}_{\ell}=\phi^{j^{\prime}}_{j_{\ell}-1;x_{j_{\ell}}}(\bm{x}^{\prime})$. Then we have
@@
Similar to Remark 2, for any $j,j^{\prime}\in[n]$ such that $j\leq j^{\prime}$ and $d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=1$, then by Observation 2, we can let $\big{(}S\cap[1,j-1]\big{)}\cup\big{(}T^{L}\cap[j+1,j^{\prime}]\big{)}\cup\big{(}S\cap[j^{\prime}+1,n]\big{)}=\{j_{1}\}$. If $j_{1}\in[1,j-1]\cup[j^{\prime}+1,n]$, then we have
$B^{\text{S}}_{1}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})$ =$=\big{\{}\phi^{j}_{j_{1};a}(\bm{x}):a\in\Sigma_{q}\big{\}}$
=$=\big{\{}\phi^{j^{\prime}}_{j_{1};a}(\bm{x}^{\prime}):a\in\Sigma_{q}\big{\}};$
@@
$B^{\text{S}}_{1}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})$ =$=\big{\{}\phi^{j}_{j_{1};a}(\bm{x}):a\in\Sigma_{q}\big{\}}$
=$=\big{\{}\phi^{j^{\prime}}_{j_{1}-1;a}(\bm{x}^{\prime}):a\in\Sigma_{q}\big{\}}.$
@@ III Methodology
In this section, we will always assume that $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$ are arbitrarily chosen such that $d=d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 2$. By Observation 1, we have $B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})=\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})$, where $\Lambda=\Lambda(\bm{x},\bm{x}^{\prime})$ is defined by (II-B). To find the size of $B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})$, we will develop a method to divide the set $\Lambda$, and correspondingly the set $B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})$, into some subsets such that the size of each subset can be easily obtained.
Specifically, define some subsets as follows:
  • For each $\ell\in\{0,1,2\}$, let
    $\Lambda_{\ell}=\Lambda_{\ell}(\bm{x},\bm{x}^{\prime})$$\triangleq\left\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j,j^{\prime}\in[n]\right.$
    $~{}~{}~{}~{}~{}~{}\left.\text{and}~{}d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})=\ell\right\}$
@@
    $\Omega_{\ell}=\Omega_{\ell}(\bm{x},\bm{x}^{\prime})\triangleq\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda_{\ell}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime}).$
  • For each $\ell\in\{0,1,2\}$, let
    $\Lambda^{L}_{\ell}=\Lambda^{L}_{\ell}(\bm{x},\bm{x}^{\prime})$$\triangleq\left\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):(j,j^{\prime})\in[n]\times[n],\right.$
    $~{}~{}~{}~{}~{}\left.j\leq j^{\prime}~{}\text{and}~{}d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=\ell\right\}$
@@
    $\Omega^{L}_{\ell}=\Omega^{L}_{\ell}(\bm{x},\bm{x}^{\prime})\triangleq\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda^{L}_{\ell}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime}).$
  • For each $\ell\in\{0,1,2\}$, let
    $\Lambda^{R}_{\ell}=\Lambda^{R}_{\ell}(\bm{x},\bm{x}^{\prime})$$\triangleq\left\{(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j}):(j,j^{\prime})\in[n]\times[n],\right.$
    $~{}~{}~{}~{}~{}\left.j\leq j^{\prime}~{}\text{and}~{}d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})=\ell\right\}$
@@
By the above definitions, we have
$\Lambda_{\ell}=\Lambda^{L}_{\ell}\cup\Lambda^{R}_{\ell}$
for each $\ell\in\{0,1,2\}$ and
@@
Correspondingly, we have
$\Omega_{\ell}=\Omega^{L}_{\ell}\cup\Omega^{R}_{\ell}$
for each $\ell\in\{0,1,2\}$ and
@@
We remark that $(\bm{x},\bm{x}^{\prime})$ should be viewed as an ordered pair in the notations $\Lambda^{X}_{\ell}(\bm{x},\bm{x}^{\prime})$, where $X\in\{L,R\}$ and $\ell\in\{0,1,2\}$. Moreover, by the definitions, $(\bm{z},\bm{z}^{\prime})\in\Lambda^{R}_{\ell}=\Lambda^{R}_{\ell}(\bm{x},\bm{x}^{\prime})$ if and only if $(\bm{z}^{\prime},\bm{z})\in\Lambda^{L}_{\ell}(\bm{x}^{\prime},\bm{x})$.
@@ III-A For $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})$ and $\Lambda^{R}_{0}(\bm{x},\bm{x}^{\prime})$
We first consider $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})$. Let $S$ and $T^{L}$ be defined according to (3) and (4), respectively.
@@
    If $|T^{L}\cap[i_{1}+1,i_{d}]|=0$, then $x_{[n]\backslash i_{1}}=x^{\prime}_{[n]\backslash i_{d}}$ and $\Lambda^{L}_{0}=\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$. Hence, we have $\Omega^{L}_{0}=\Omega^{L}_{0}(\bm{x},\bm{x}^{\prime})=B^{\text{S}}_{1}(x_{[n]\backslash i_{1}})=B^{\text{S}}_{1}(x^{\prime}_{[n]\backslash i_{d}})$.
@@
Note that $j\leq j^{\prime}$ and $S=\{i_{1},\cdots,i_{d}\}$ such that $i_{1}<\cdots<i_{d}$. Then from the conditions $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=0$, we have $j\leq i_{1}$ and $j^{\prime}\geq i_{d}$, which implies $T^{L}\cap[i_{1}+1,i_{d}]\subseteq T^{L}\cap[j+1,j^{\prime}]$. Further, from the condition $|T^{L}\cap[j+1,j^{\prime}]|=0$, we have $|T^{L}\cap[i_{1}+1,i_{d}]|\leq|T^{L}\cap[j+1,j^{\prime}]|=0$. Thus, if $|T^{L}\cap[i_{1}+1,i_{d}]|\geq 1$, then there is no $(j,j^{\prime})$ that satisfies the conditions in (6), and so $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})=\emptyset$.
Conversely, if $|T^{L}\cap[i_{1}+1,i_{d}]|=0$, then by the definition of $k_{a},k^{\prime}_{a},T^{L}$ and $S$, it is not hard to see that $x_{[n]\backslash i_{1}}=x^{\prime}_{[n]\backslash i_{d}}$ and $(j,j^{\prime})$ satisfies (6) if and only if $k_{a}\leq j\leq i_{1}<i_{d}\leq j^{\prime}\leq k^{\prime}_{a}$. Therefore, we have $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{a}\leq j\leq i_{1}<i_{d}\leq j^{\prime}\leq k^{\prime}_{a}\}\neq\emptyset.$ Moreover, by the definition of $k_{a}$ and $k^{\prime}_{a}$, we can obtain $[k_{a},i_{1}-1]\cap S=\emptyset$ and $[k_{a}+1,i_{1}]\cap T^{L}=\emptyset$, so by 1) of Lemma 1, $x_{[k_{a},i_{1}]}$ is contained in a run of $\bm{x}$. Similarly, we can obtain $[i_{d}+1,k^{\prime}_{a}]\cap S=[i_{d}+1,k^{\prime}_{a}]\cap T^{L}=\emptyset$, and so by 2) of Lemma 1, $x^{\prime}_{[i_{d},k^{\prime}_{a}]}$ is contained in a run of $\bm{x}^{\prime}$. Thus, we have $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{a}\leq j\leq i_{1}<i_{d}\leq j^{\prime}\leq k^{\prime}_{a}\}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$.
@@
    If $|T^{R}\cap[i_{1}+1,i_{d}]|=0$, then $x_{[n]\backslash i_{d}}=x^{\prime}_{[n]\backslash i_{1}}$ and $\Lambda^{R}_{0}=\Lambda^{R}_{0}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})\}$. Hence, we have $\Omega^{R}_{0}=\Omega^{R}_{0}(\bm{x},\bm{x}^{\prime})=B^{\text{S}}_{1}(x_{[n]\backslash i_{d}})=B^{\text{S}}_{1}(x^{\prime}_{[n]\backslash i_{1}})$.

### v2

@@ II-C The notation $\phi$
Suppose $\bm{x}=01010111$ and $\bm{x}^{\prime}=01101011$. Then we have $B^{\text{S}}_{1}(x_{[n]\backslash 4},x^{\prime}_{[n]\backslash 7})=\{\phi^{4}_{7;x^{\prime}_{6}}(\bm{x}),\phi^{4}_{3;x^{\prime}_{3}}(\bm{x})\}=\{\phi^{7}_{3;x_{3}}(\bm{x}^{\prime}),\phi^{7}_{6;x_{7}}(\bm{x}^{\prime})\}$. In fact, we can easily check that $x_{[n]\backslash 4}=0100111$ and $\bm{x}^{\prime}_{[n]\backslash 7}=0110101$. Moreover, we can find $B^{\text{S}}_{1}(x_{[n]\backslash 4},x^{\prime}_{[n]\backslash 7})=\{0100101,0110111\}$, $\phi^{4}_{7;x^{\prime}_{6}}(\bm{x})=0100101=\phi^{7}_{3;x_{3}}(\bm{x}^{\prime})$ and $\phi^{4}_{3;x^{\prime}_{3}}(\bm{x})=0110111=\phi^{7}_{6;x_{7}}(\bm{x}^{\prime})$.
@@
For any $j,j^{\prime}\in[n]$ such that $j\leq j^{\prime}$, if $d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=2$, then by Observation 2, we can denote $\{j_{1},j_{2}\}=\big{(}S\cap[1,j-1]\big{)}\cup\big{(}T^{L}\cap[j+1,j^{\prime}]\big{)}\cup\big{(}S\cap[j^{\prime}+1,n]\big{)}$. For each $\ell\in\{1,2\}$: if $j_{\ell}\in[1,j-1]\cup[j^{\prime}+1,n]$, let $\bm{z}_{\ell}=\phi^{j}_{j_{\ell};x^{\prime}_{j_{\ell}}}(\bm{x})$ and $\bm{w}_{\ell}=\phi^{j^{\prime}}_{j_{\ell};x_{j_{\ell}}}(\bm{x}^{\prime})$; if $j_{\ell}\in[j+1,j^{\prime}]$, let $\bm{z}_{\ell}=\phi^{j}_{j_{\ell};x^{\prime}_{j_{\ell}-1}}(\bm{x})$ and $\bm{w}_{\ell}=\phi^{j^{\prime}}_{j_{\ell}-1;x_{j_{\ell}}}(\bm{x}^{\prime})$. Then we have
@@
Similar to Remark 2, for any $j,j^{\prime}\in[n]$ such that $j\leq j^{\prime}$ and $d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=1$, then by Observation 2, we can denote $\{j_{1}\}=\big{(}S\cap[1,j-1]\big{)}\cup\big{(}T^{L}\cap[j+1,j^{\prime}]\big{)}\cup\big{(}S\cap[j^{\prime}+1,n]\big{)}$. If $j_{1}\in[1,j-1]\cup[j^{\prime}+1,n]$, then we have
$B^{\text{S}}_{1}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=\big{\{}\phi^{j}_{j_{1};a}(\bm{x}):a\in\Sigma_{q}\big{\}}=\big{\{}\phi^{j^{\prime}}_{j_{1};a}(\bm{x}^{\prime}):a\in\Sigma_{q}\big{\}};$
@@
$B^{\text{S}}_{1}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=\big{\{}\phi^{j}_{j_{1};a}(\bm{x}):a\in\Sigma_{q}\big{\}}=\big{\{}\phi^{j^{\prime}}_{j_{1}-1;a}(\bm{x}^{\prime}):a\in\Sigma_{q}\big{\}}.$
@@ III Methodology
In this section, we will always assume that $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$ are arbitrarily chosen such that $d=d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 2$. By Observation 1, we have $B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})=\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})$, where $\Lambda=\Lambda(\bm{x},\bm{x}^{\prime})$ is defined by (2). To find the size of $B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})$, we will develop a method to divide the set $\Lambda$, and correspondingly the set $B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})$, into some subsets that can be easily obtained from $\bm{x}$ and $\bm{x}^{\prime}$. See Fig. 1 for an overview of the dividing of the set $\Lambda$.
Figure 1: An overview of the dividing of $\Lambda$, where $\Lambda=\big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j,j^{\prime}\in[n]~{}\text{and}~{}d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})\leq 2\big{\}}$ is defined by (2). First, $\Lambda$ is divided into $\Lambda_{0},\Lambda_{1}$ and $\Lambda_{2}$ according to the value of $d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})$. Then for each $\ell\in\{0,1,2\}$, $\Lambda_{\ell}$ is divided into $\Lambda_{\ell}^{L}$ and $\Lambda_{\ell}^{R}$ according to the relationship of $j$ and $j^{\prime}$. Here we assume $j\leq j^{\prime}$ and consider $(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})$ and $(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})$. Finally, for each $\ell\in\{1,2\}$ and each $X\in\{L,R\}$, $\Lambda_{\ell}^{X}$ is divided into $\Lambda_{\ell,i}^{X}$, $i=1,\cdots,p_{\ell}$, where $p_{1}=3$ and $p_{2}=6$, according to the value of $(|S\cap[1,j-1]|,|T^{X}\cap[j+1,j^{\prime}]|,|S\cap[j^{\prime}+1,n]|)$, where by Observation 2, $d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})=|S\cap[1,j-1]|+|T^{X}\cap[j+1,j^{\prime}]|+|S\cap[j^{\prime}+1,n]|$. Moreover, the sets $\Lambda^{X}_{0}$ and $\Lambda^{X}_{\ell,i}$ can be easily obtained from $\bm{x}$ and $\bm{x}^{\prime}$.
Definition 1
For each $\ell\in\{0,1,2\}$:
  • let
    $\Lambda_{\ell}=\Lambda_{\ell}(\bm{x},\bm{x}^{\prime})\triangleq\left\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j,j^{\prime}\in[n]~{}\text{and}~{}d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})=\ell\right\}$
@@
    $\Omega_{\ell}=\Omega_{\ell}(\bm{x},\bm{x}^{\prime})\triangleq\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda_{\ell}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime});$
  • let
    $\Lambda^{L}_{\ell}=\Lambda^{L}_{\ell}(\bm{x},\bm{x}^{\prime})$$\triangleq\left\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):(j,j^{\prime})\in[n]\times[n],j\leq j^{\prime}~{}\text{and}~{}d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=\ell\right\}$
@@
    $\Omega^{L}_{\ell}=\Omega^{L}_{\ell}(\bm{x},\bm{x}^{\prime})\triangleq\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda^{L}_{\ell}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime});$
  • let
    $\Lambda^{R}_{\ell}=\Lambda^{R}_{\ell}(\bm{x},\bm{x}^{\prime})$$\triangleq\left\{(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j}):(j,j^{\prime})\in[n]\times[n],j\leq j^{\prime}~{}\text{and}~{}d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})=\ell\right\}$
@@
By the above definitions, we have $\Lambda_{\ell}=\Lambda^{L}_{\ell}\cup\Lambda^{R}_{\ell}$ for each $\ell\in\{0,1,2\}$ and
@@
Correspondingly, we have $\Omega_{\ell}=\Omega^{L}_{\ell}\cup\Omega^{R}_{\ell}$ for each $\ell\in\{0,1,2\}$ and
@@
We remark that $(\bm{x},\bm{x}^{\prime})$ should be viewed as an ordered pair in the notations $\Lambda^{X}_{\ell}(\bm{x},\bm{x}^{\prime})$, $X\in\{L,R\}$ and $\ell\in\{0,1,2\}$. By the definitions, $(\bm{z},\bm{z}^{\prime})\in\Lambda^{R}_{\ell}=\Lambda^{R}_{\ell}(\bm{x},\bm{x}^{\prime})$ if and only if $(\bm{z}^{\prime},\bm{z})\in\Lambda^{L}_{\ell}(\bm{x}^{\prime},\bm{x})$.
@@ III-A For $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})$ and $\Lambda^{R}_{0}(\bm{x},\bm{x}^{\prime})$
We first consider $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})$. Let $S$ and $T^{L}$ be defined by (3) and (4), respectively. We have the following claim.
@@
    If $|T^{L}\cap[i_{1}+1,i_{d}]|=0$, then $x_{[n]\backslash i_{1}}=x^{\prime}_{[n]\backslash i_{d}}$ and $\Lambda^{L}_{0}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$. Hence, we have $\Omega^{L}_{0}=B^{\text{S}}_{1}(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})=B^{\text{S}}_{1}(x_{[n]\backslash i_{1}})=B^{\text{S}}_{1}(x^{\prime}_{[n]\backslash i_{d}})$.
@@
Note that $j\leq j^{\prime}$ and $S=\{i_{1},\cdots,i_{d}\}$ such that $i_{1}<\cdots<i_{d}$. Then from the conditions $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=0$, we have $j\leq i_{1}$ and $j^{\prime}\geq i_{d}$, which implies
$T^{L}\cap[i_{1}+1,i_{d}]\subseteq T^{L}\cap[j+1,j^{\prime}].$
Combining this with the condition $|T^{L}\cap[j+1,j^{\prime}]|=0$, we have
$|T^{L}\cap[i_{1}+1,i_{d}]|\leq|T^{L}\cap[j+1,j^{\prime}]|=0.$
Thus, if $|T^{L}\cap[i_{1}+1,i_{d}]|\geq 1$, then there is no $(j,j^{\prime})$ that satisfies (6), and so $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})=\emptyset$.
Conversely, if $|T^{L}\cap[i_{1}+1,i_{d}]|=0$, then clearly, we have $x_{[n]\backslash i_{1}}=x^{\prime}_{[n]\backslash i_{d}}$. Moreover, by the definition of $k_{a},k^{\prime}_{a},T^{L}$ and $S$, it is not hard to see that $(j,j^{\prime})$ satisfies (6) if and only if $k_{a}\leq j\leq i_{1}<i_{d}\leq j^{\prime}\leq k^{\prime}_{a}$ (see Fig. 2). Therefore, we have $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{a}\leq j\leq i_{1}<i_{d}\leq j^{\prime}\leq k^{\prime}_{a}\}\neq\emptyset.$ Moreover, by the definition of $k_{a}$ and $k^{\prime}_{a}$, we can obtain $[k_{a},i_{1}-1]\cap S=\emptyset$ and $[k_{a}+1,i_{1}]\cap T^{L}=\emptyset$, so by 1) of Lemma 1, $x_{[k_{a},i_{1}]}$ is contained in a run of $\bm{x}$. Similarly, we can obtain $[i_{d}+1,k^{\prime}_{a}]\cap S=[i_{d}+1,k^{\prime}_{a}]\cap T^{L}=\emptyset$, and so by 2) of Lemma 1, $x^{\prime}_{[i_{d},k^{\prime}_{a}]}$ is contained in a run of $\bm{x}^{\prime}$. Thus, we have $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{a}\leq j\leq i_{1}<i_{d}\leq j^{\prime}\leq k^{\prime}_{a}\}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$.
Figure 2: An illustration of the pair $(j,j^{\prime})$ satisfying (6). Each black dot represents a symbol of $\bm{x}$ (in the upper row) or a symbol of $\bm{x}^{\prime}$ (in the lower row). Symbols are connected by a solid segment are identical, while those connected by a dashed segment are distinct. Here, $k_{a}=\max(T^{L}\cap[1,i_{1}])$ and $k^{\prime}_{a}=n$ because $T^{L}\cap[i_{d}+1,n]=\emptyset$. We can find that $(j,j^{\prime})$ satisfies (6) if and only if $k_{a}\leq j\leq i_{1}<i_{d}\leq j^{\prime}\leq k^{\prime}_{a}$. In this example, $d=4$. Moreover, we can find that $x_{i}=x^{\prime}_{i}=x_{i+1}$ for each $i\in[k_{a},i_{1}-1]$ and $x^{\prime}_{i}=x_{i}=x^{\prime}_{i-1}$ for each $i\in[i_{d}+1,k_{a}^{\prime}]$. Hence, $x_{[k_{a},i_{1}]}$ is contained in a run of $\bm{x}$ and $x^{\prime}_{[i_{d},k^{\prime}_{a}]}$ is contained in a run of $\bm{x}^{\prime}$.
@@
    If $|T^{R}\cap[i_{1}+1,i_{d}]|=0$, then $x_{[n]\backslash i_{d}}=x^{\prime}_{[n]\backslash i_{1}}$ and $\Lambda^{R}_{0}=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})\}$. Hence, we have $\Omega^{R}_{0}=B^{\text{S}}_{1}(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})=B^{\text{S}}_{1}(x_{[n]\backslash i_{d}})=B^{\text{S}}_{1}(x^{\prime}_{[n]\backslash i_{1}})$.